What is the physical significance of the factor(9/5)
in Equation T(F)=(9/5)T(C)+32? Why is this factor missing
in Equation T(C)=T(K)-273?

Grade:11

1 Answers

Nirmal Singh.
10 years ago
Because In degree Celcius ice point is O degree Celcius and Steam point is 100 Degree Celcius.and In Kelvin ice point is 273K and Steam point is 373 K and in Fahrenheit the ice point is 32 F and Steam point is 212 F. So we can see the A change of 1F means 1/180 of the interval between the steam and the ice point and a change of 1 C & 1 K means 1/100 of the interval between the steam and the ice point. Due this reason the factor 9 / 5 = 180 /100 in T(F)=(9/5)T(C)+32 and not in T(C)=T(K)-273 because 100 /100.
Regards,
Nirmal SIngh
Askiitians Faculty